# Accepted to 2024 European Conference on Computer Vision

### v2

@@ III-B2 Threshold Retrospective Gradient (TRG)
Biological systems exhibit a negative correlation with the preceding rate of depolarization, which indicates the excitatory status. We propose considering a historical rate of change for the threshold $\Theta(t)$. This is because focusing solely on the preceding rate of depolarization as implemented in [30] is insufficient and can result in an unstable or overly sensitive dynamic threshold. This approach lacks memory, reacting too quickly to transient changes without considering longer-term neuronal activity patterns. Consequently, this can result in excessive threshold fluctuations and potentially amplify noise, undermining network stability and reliability. Therefore, we take into account the average of the historical rate of change of the dynamic threshold. Utilizing the historical rate of change of the dynamic threshold to calculate the current threshold can provide the neuron with inertia or memory of its past states. This can be useful in situations where the neuron should react not only based on the current state but also considering its past behavior. In discrete terms, if $\Theta(t)$ is the threshold at time $t$, then the rate of change at $t$ is:
@@ IV-A Experimental Protocol
We assess the effectiveness of the proposed ABN in two Computer Vision tasks: Image Classification for the N-MNIST [9], CIFAR10-DVS [13], N-ImageNet [12], DVS128 gesture recognition [11] datasets, and Object Semantic Segmentation for the ESD-1 [10] and ESD-2 [10] datasets. The details of the datasets are explained in supplementary section B. The Spiking MLP employs Spatio-Temporal Backpropagation [30] for training, allowing the network to consider both spatial and temporal information during the learning process. Although the input to the first layer should be a spike train, we can directly use asynchronous event-based vision data without requiring specific conversion. The initialization of network parameters, such as weights and thresholds, is crucial for stabilizing the network’s firing activities. We adopt the initialization strategy as discussed in [37]. To make a fair comparison with [30], we have used a similar experimental configuration. The model was trained on a GPU PC featuring 128GB of memory, an Intel Xeon W-2155 Processor, and NVIDIA Quadro RTX 8000 48GB graphics cards, housed in a Lenovo ThinkStation P520. While this is essential for assessing the efficacy of the proposed method, implementing it directly on neuromorphic hardware may require additional engineering efforts, which are beyond the scope of this work. Custom code was written to develop the LIF and SRM within this framework. Synaptic weights are initialized randomly, and a decay constant, $\tau_{1}$, is set to 0.1 ms. The weights for constants $k_{1}$, $k_{2}$, and $k_{3}$ are experimentally selected as 0.25, 0.50, and 0.25, respectively (see supplementary document).
@@ IV-B Comparative Results across Diverse Datasets
Method N-MNIST DVS128 CIFAR10 SRM LIF SRM LIF SRM LIF Spiking RBM [38] 92.1 93.16 87.38 90.21 83.04 86.25 Spiking MLP (BP) [39] 94.52 97.66 90.72 93.4 86.23 90.06 Spiking MLP (STDP) [40] 93.47 95 90.8 92.01 87.5 91.7 Spiking MLP (STBP) [37] 97.13 98.89 92.54 93.64 86.23 91.73 DT1 [27] 99.05 99.4 95.01 96.88 89.17 92.65 DT2 [29] 98.13 98.24 92.54 95.54 89.38 91.47 BDETT [30] 99.15 99.45 94.09 96.05 91.61 93.5 ABN (Ours) 99.23 99.48 95.64 98.74 93.5 94.74
@@
Method ESD-1 ESD-2 SRM LIF SRM LIF Spiking RBM [38] 48.95 51.05 45.2 50.31 Spiking MLP (BP) [39] 49.53 54 45.5 52.42 Spiking MLP (STDP) [40] 52.63 58.84 46.7 49.44 Spiking MLP (STBP) [37] 55.82 63.09 49.07 51.37 DT1 [27] 57.55 61.45 50.13 54.59 DT2 [29] 58.28 64.3 53.49 56.5 BDETT [30] 61.02 65.39 51.36 55.46 ABN (Ours) 61.56 67.29 56.23 58.04
@@ IV-C Evaluation on Degraded Input on ESD-1
Exp 1: varying clutter objects, Bright light, 62cm height, Rotational motion, 0.15 m/s speed Method 2 Obj 4 Obj 6 Obj 8 Obj 10 Obj Spiking MLP (STBP) [37] 74.16% 73.05% 66.79% 66.42% 54.42% DT1 [27] 77.56% 74.65% 68.20% 68.70% 55.49% DT2 [29] 78.46% 75.42% 69.28% 69.62% 56.16% BDETT [30] 79.07% 78.03% 74.14% 69.83% 58.28% ABN (ours) 81.49% 79.74% 74.97% 72.16% 60.03% Exp 2: 6 Objects, varying lighting conditions, 62cm height, Rotational Motion, 0.15 m/s speed. Method Bright Light Low light Spiking MLP (STBP) [37] 58.73% 61.59% DT1 [27] 61.32% 62.91% DT2 [29] 61.24% 64.53% BDETT [30] 63.92% 66.03% ABN (Ours) 65.98% 67.42% Exp 3: 6 Objects, Bright Light, 62cm height, Varying directions of motion, 0.15 m/s speed. Method Linear Rotational Partial Rotational Spiking MLP (STBP) [37] 48.09% 63.99% 67.43% DT1 [27] 49.09% 65.14% 69.31% DT2 [29] 50.43% 66.89% 69.79% BDETT [30] 52.09% 68.60% 72.04% ABN (Ours) 55.50% 69.45% 73.54% Exp 4: 6 Objects, Bright Light, 62cm height, Rotational motion, Varying speed. Method 0.15 m/s 0.3 m/s 0.1 m/s Spiking MLP (STBP) [37] 53.16% 68.12% 56.29% DT1 [27] 54.56% 70.27% 60.16% DT2 [29] 55.27% 72.31% 59.63% BDETT [30] 56.67% 75.09% 63.16% ABN (Ours) 59.63% 76.72% 63.56% Exp 5: 6 Objects, Bright Light, Varying camera height, Rotational motion, Varying speed. Method 62 cm 82 cm Spiking MLP (STBP) [37] 62.93% 55.05% DT1 [27] 63.82% 59.08% DT2 [29] 65.63% 59.93% BDETT [30] 65.82% 59.00% ABN (ours) 68.40% 63.27%
@@ IV-D Homeostasis Evaluation
$FR\_m$ $\Delta FR\_m$ $FR\_m\_std$ $\Delta FR\_m\_std$ $FR\_s\_std$ $\Delta FR\_s\_std$ Ideal Condition Spiking MLP (STBP) [37] 0.631 0.485 0.001564 DT1 [27] 0.524 0.459 0.001146 DT2 [29] 0.507 0.416 0.000983 BDETT [30] 0.436 0.403 0.001224 ABN (Ours) 0.385 0.253 0.000891 Occlusion Spiking MLP (STBP) [37] 0.783 0.152 0.582 0.097 0.001396 0.000168 DT1 [27] 0.692 0.168 0.561 0.102 0.003963 0.002817 DT2 [29] 0.586 0.079 0.484 0.068 0.002724 0.001741 BDETT [30] 0.47 0.034 0.357 0.046 0.002291 0.001067 ABN (Ours) 0.373 0.012 0.291 0.038 0.000784 0.000107 Light Spiking MLP (STBP)[37] 0.786 0.155 0.594 0.109 0.003942 0.002378 DT1 [27] 0.698 0.174 0.405 0.054 0.002508 0.001362 DT2 [29] 0.594 0.087 0.489 0.073 0.003641 0.002658 BDETT [30] 0.498 0.062 0.467 0.064 0.001157 0.000067 ABN (Ours) 0.405 0.02 0.295 0.042 0.000889 0.000002 Speed Spiking MLP (STBP) [37] 0.746 0.115 0.582 0.097 0.001666 0.000102 DT1 [27] 0.62 0.096 0.406 0.053 0.001167 0.000021 DT2 [29] 0.609 0.102 0.351 0.065 0.002839 0.001856 BDETT [30] 0.517 0.081 0.352 0.051 0.002077 0.000853 ABN (Ours) 0.305 0.08 0.206 0.047 0.000881 0.00001 Motion Direction Spiking MLP (STBP) [37] 0.701 0.07 0.505 0.02 0.003798 0.002234 DT1 [27] 0.601 0.077 0.409 0.05 0.003461 0.002315 DT2 [29] 0.415 0.092 0.397 0.019 0.002827 0.001844 BDETT[30] 0.395 0.041 0.351 0.052 0.00133 0.000106 ABN (Ours) 0.378 0.007 0.243 0.01 0.000875 0.000016 Size Variance Spiking MLP (STBP) [37] 0.759 0.128 0.533 0.048 0.003901 0.002337 DT1 [27] 0.675 0.151 0.449 0.01 0.002263 0.001117 DT2 [29] 0.551 0.044 0.45 0.034 0.001346 0.000363 BDETT [30] 0.481 0.045 0.354 0.049 0.002763 0.001539 ABN (Ours) 0.373 0.012 0.255 0.002 0.000879 0.000012
In an experimental setup across all variations of the ESD-1 dataset, various statistical indicators were used to quantify the homeostasis of the host SNNs. Specifically, for a fair comparison with [30] we employed $FR_{m}$, $FR_{m\_std}$, and $FR_{s\_std}$ metrics, representing the average neuron firing rate over all trials, the mean of standard deviations of neuron firing rates over individual trials, and the variability of these standard deviations, respectively. Our proposed ABN model demonstrates significantly lower values across all three metrics and all conditions. These results suggest a stable and well-regulated network, highlighting the ABN model’s capacity for homeostasis.
@@ IV-E Energy Consumption
Table V presents an analysis of power consumption, comparing CNN, SotA SNNs and the ABN using the DVS128 gesture recognition dataset. As proposed in [31], MAC (Multiply-Accumulate operations) and AC (Accumulate Count) were used to estimate power consumption. A detailed description of power consumption calculations is provided in the supplementary document.
CNN models such as [41] and [42], show a higher power consumption of 0.541W and 1.56W, respectively, despite their zero AC. In contrast, using SNN models leads to significantly lower power usage, with hybrid and SNN models [31] consuming 0.404W and 0.053W, respectively. BDETT’s SNN model [31] further reduces this consumption to 0.046W. Notably, our proposed SNN model outperforms these by achieving the lowest power consumption of just 0.038W, even with a higher AC of 105, underscoring the efficiency of our model in terms of energy usage compared to both CNNs and other SNNs.
@@
Method Model MAC AC Power (W) Calabrese [41] CNN 285 0 0.541 Baldwin [42] CNN 963 0 1.56 Asude [31] Hybrid 235 81 0.404 Asude [31] SNN 0.6 123 0.053 BDETT [30] SNN 0.7 115 0.046 Ours SNN 0.4 105 0.038

### v3

@@ III-B2 Threshold Retrospective Gradient (TRG)
Biological systems exhibit a negative correlation with the preceding rate of depolarization, which indicates the excitatory status. We propose considering a historical rate of change for the threshold $\Theta(t)$. This is because focusing solely on the preceding rate of depolarization as implemented in [Ding2022BiologicallyNetworks] is insufficient and can result in an unstable or overly sensitive dynamic threshold. This approach lacks memory, reacting too quickly to transient changes without considering longer-term neuronal activity patterns. Consequently, this can result in excessive threshold fluctuations and potentially amplify noise, undermining network stability and reliability. Therefore, we take into account the average of the historical rate of change of the dynamic threshold. Utilizing the historical rate of change of the dynamic threshold to calculate the current threshold can provide the neuron with inertia or memory of its past states. This can be useful in situations where the neuron should react not only based on the current state but also considering its past behavior. In discrete terms, if $\Theta(t)$ is the threshold at time $t$, then the rate of change at $t$ is:
@@ IV-A Experimental Protocol
We assess the effectiveness of the proposed ABN in two Computer Vision tasks: Image Classification for the N-MNIST [Orchard2015ConvertingSaccades], CIFAR10-DVS [Li2017CIFAR10-DVS:Classification], N-ImageNet [kim2021n], DVS128 gesture recognition [Amir2017ASystem] datasets, and Object Semantic Segmentation for the ESD-1 [Huang2023AEnvironment] and ESD-2 [Huang2023AEnvironment] datasets. The details of the datasets are explained in supplementary section B. The Spiking MLP employs Spatio-Temporal Backpropagation [Ding2022BiologicallyNetworks] for training, allowing the network to consider both spatial and temporal information during the learning process. Although the input to the first layer should be a spike train, we can directly use asynchronous event-based vision data without requiring specific conversion. The initialization of network parameters, such as weights and thresholds, is crucial for stabilizing the network’s firing activities. We adopt the initialization strategy as discussed in [Wu2018Spatio-TemporalNetworks]. To make a fair comparison with [Ding2022BiologicallyNetworks], we have used a similar experimental configuration. The model was trained on a GPU PC featuring 128GB of memory, an Intel Xeon W-2155 Processor, and NVIDIA Quadro RTX 8000 48GB graphics cards, housed in a Lenovo ThinkStation P520. While this is essential for assessing the efficacy of the proposed method, implementing it directly on neuromorphic hardware may require additional engineering efforts, which are beyond the scope of this work. Custom code was written to develop the LIF and SRM within this framework. Synaptic weights are initialized randomly, and a decay constant, $\tau_{1}$, is set to 0.1 ms. The weights for constants $k_{1}$, $k_{2}$, and $k_{3}$ are experimentally selected as 0.25, 0.50, and 0.25, respectively (see supplementary document).
@@ IV-B Comparative Results across Diverse Datasets
Method N-MNIST DVS128 CIFAR10 SRM LIF SRM LIF SRM LIF Spiking RBM [Neftci2014Event-drivenSystems] 92.1 93.16 87.38 90.21 83.04 86.25 Spiking MLP (BP) [OConnor2016DeepNetworks] 94.52 97.66 90.72 93.4 86.23 90.06 Spiking MLP (STDP) [Ferre2018UnsupervisedSTDP] 93.47 95 90.8 92.01 87.5 91.7 Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 97.13 98.89 92.54 93.64 86.23 91.73 DT1 [Hao2018ARuleb] 99.05 99.4 95.01 96.88 89.17 92.65 DT2 [Kim2021SpikingUpdate.] 98.13 98.24 92.54 95.54 89.38 91.47 BDETT [Ding2022BiologicallyNetworks] 99.15 99.45 94.09 96.05 91.61 93.5 ABN (Ours) 99.23 99.48 95.64 98.74 93.5 94.74
@@
Method ESD-1 ESD-2 SRM LIF SRM LIF Spiking RBM [Neftci2014Event-drivenSystems] 48.95 51.05 45.2 50.31 Spiking MLP (BP) [OConnor2016DeepNetworks] 49.53 54 45.5 52.42 Spiking MLP (STDP) [Ferre2018UnsupervisedSTDP] 52.63 58.84 46.7 49.44 Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 55.82 63.09 49.07 51.37 DT1 [Hao2018ARuleb] 57.55 61.45 50.13 54.59 DT2 [Kim2021SpikingUpdate.] 58.28 64.3 53.49 56.5 BDETT [Ding2022BiologicallyNetworks] 61.02 65.39 51.36 55.46 ABN (Ours) 61.56 67.29 56.23 58.04
@@ IV-C Evaluation on Degraded Input on ESD-1
Exp 1: varying clutter objects, Bright light, 62cm height, Rotational motion, 0.15 m/s speed Method 2 Obj 4 Obj 6 Obj 8 Obj 10 Obj Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 74.16% 73.05% 66.79% 66.42% 54.42% DT1 [Hao2018ARuleb] 77.56% 74.65% 68.20% 68.70% 55.49% DT2 [Kim2021SpikingUpdate.] 78.46% 75.42% 69.28% 69.62% 56.16% BDETT [Ding2022BiologicallyNetworks] 79.07% 78.03% 74.14% 69.83% 58.28% ABN (ours) 81.49% 79.74% 74.97% 72.16% 60.03% Exp 2: 6 Objects, varying lighting conditions, 62cm height, Rotational Motion, 0.15 m/s speed. Method Bright Light Low light Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 58.73% 61.59% DT1 [Hao2018ARuleb] 61.32% 62.91% DT2 [Kim2021SpikingUpdate.] 61.24% 64.53% BDETT [Ding2022BiologicallyNetworks] 63.92% 66.03% ABN (Ours) 65.98% 67.42% Exp 3: 6 Objects, Bright Light, 62cm height, Varying directions of motion, 0.15 m/s speed. Method Linear Rotational Partial Rotational Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 48.09% 63.99% 67.43% DT1 [Hao2018ARuleb] 49.09% 65.14% 69.31% DT2 [Kim2021SpikingUpdate.] 50.43% 66.89% 69.79% BDETT [Ding2022BiologicallyNetworks] 52.09% 68.60% 72.04% ABN (Ours) 55.50% 69.45% 73.54% Exp 4: 6 Objects, Bright Light, 62cm height, Rotational motion, Varying speed. Method 0.15 m/s 0.3 m/s 0.1 m/s Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 53.16% 68.12% 56.29% DT1 [Hao2018ARuleb] 54.56% 70.27% 60.16% DT2 [Kim2021SpikingUpdate.] 55.27% 72.31% 59.63% BDETT [Ding2022BiologicallyNetworks] 56.67% 75.09% 63.16% ABN (Ours) 59.63% 76.72% 63.56% Exp 5: 6 Objects, Bright Light, Varying camera height, Rotational motion, Varying speed. Method 62 cm 82 cm Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 62.93% 55.05% DT1 [Hao2018ARuleb] 63.82% 59.08% DT2 [Kim2021SpikingUpdate.] 65.63% 59.93% BDETT [Ding2022BiologicallyNetworks] 65.82% 59.00% ABN (ours) 68.40% 63.27%
@@ IV-D Homeostasis Evaluation
FR_mFR\_m $\Delta FR\_m$ FR_m_stdFR\_m\_std $\Delta FR\_m\_std$ FR_s_stdFR\_s\_std $\Delta FR\_s\_std$ Ideal Condition Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 0.631 0.485 0.001564 DT1 [Hao2018ARuleb] 0.524 0.459 0.001146 DT2 [Kim2021SpikingUpdate.] 0.507 0.416 0.000983 BDETT [Ding2022BiologicallyNetworks] 0.436 0.403 0.001224 ABN (Ours) 0.385 0.253 0.000891 Occlusion Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 0.783 0.152 0.582 0.097 0.001396 0.000168 DT1 [Hao2018ARuleb] 0.692 0.168 0.561 0.102 0.003963 0.002817 DT2 [Kim2021SpikingUpdate.] 0.586 0.079 0.484 0.068 0.002724 0.001741 BDETT [Ding2022BiologicallyNetworks] 0.47 0.034 0.357 0.046 0.002291 0.001067 ABN (Ours) 0.373 0.012 0.291 0.038 0.000784 0.000107 Light Spiking MLP (STBP)[Wu2018Spatio-TemporalNetworks] 0.786 0.155 0.594 0.109 0.003942 0.002378 DT1 [Hao2018ARuleb] 0.698 0.174 0.405 0.054 0.002508 0.001362 DT2 [Kim2021SpikingUpdate.] 0.594 0.087 0.489 0.073 0.003641 0.002658 BDETT [Ding2022BiologicallyNetworks] 0.498 0.062 0.467 0.064 0.001157 0.000067 ABN (Ours) 0.405 0.02 0.295 0.042 0.000889 0.000002 Speed Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 0.746 0.115 0.582 0.097 0.001666 0.000102 DT1 [Hao2018ARuleb] 0.62 0.096 0.406 0.053 0.001167 0.000021 DT2 [Kim2021SpikingUpdate.] 0.609 0.102 0.351 0.065 0.002839 0.001856 BDETT [Ding2022BiologicallyNetworks] 0.517 0.081 0.352 0.051 0.002077 0.000853 ABN (Ours) 0.305 0.08 0.206 0.047 0.000881 0.00001 Motion Direction Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 0.701 0.07 0.505 0.02 0.003798 0.002234 DT1 [Hao2018ARuleb] 0.601 0.077 0.409 0.05 0.003461 0.002315 DT2 [Kim2021SpikingUpdate.] 0.415 0.092 0.397 0.019 0.002827 0.001844 BDETT[Ding2022BiologicallyNetworks] 0.395 0.041 0.351 0.052 0.00133 0.000106 ABN (Ours) 0.378 0.007 0.243 0.01 0.000875 0.000016 Size Variance Spiking MLP (STBP) [Wu2018Spatio-TemporalNetworks] 0.759 0.128 0.533 0.048 0.003901 0.002337 DT1 [Hao2018ARuleb] 0.675 0.151 0.449 0.01 0.002263 0.001117 DT2 [Kim2021SpikingUpdate.] 0.551 0.044 0.45 0.034 0.001346 0.000363 BDETT [Ding2022BiologicallyNetworks] 0.481 0.045 0.354 0.049 0.002763 0.001539 ABN (Ours) 0.373 0.012 0.255 0.002 0.000879 0.000012
In an experimental setup across all variations of the ESD-1 dataset, various statistical indicators were used to quantify the homeostasis of the host SNNs. Specifically, for a fair comparison with [Ding2022BiologicallyNetworks] we employed $FR_{m}$, $FR_{m\_std}$, and $FR_{s\_std}$ metrics, representing the average neuron firing rate over all trials, the mean of standard deviations of neuron firing rates over individual trials, and the variability of these standard deviations, respectively. Our proposed ABN model demonstrates significantly lower values across all three metrics and all conditions. These results suggest a stable and well-regulated network, highlighting the ABN model’s capacity for homeostasis.
@@ IV-E Energy Consumption
Table V presents an analysis of power consumption, comparing CNN, SotA SNNs and the ABN using the DVS128 gesture recognition dataset. As proposed in [Aydin2023APerception], MAC (Multiply-Accumulate operations) and AC (Accumulate Count) were used to estimate power consumption. A detailed description of power consumption calculations is provided in the supplementary document.
CNN models such as [Calabrese2019DHP19:Dataset] and [Baldwin2023Time-OrderedCameras], show a higher power consumption of 0.541W and 1.56W, respectively, despite their zero AC. In contrast, using SNN models leads to significantly lower power usage, with hybrid and SNN models [Aydin2023APerception] consuming 0.404W and 0.053W, respectively. BDETT’s SNN model [Aydin2023APerception] further reduces this consumption to 0.046W. Notably, our proposed SNN model outperforms these by achieving the lowest power consumption of just 0.038W, even with a higher AC of 105, underscoring the efficiency of our model in terms of energy usage compared to both CNNs and other SNNs.
@@
Method Model MAC AC Power (W) Calabrese [Calabrese2019DHP19:Dataset] CNN 285 0 0.541 Baldwin [Baldwin2023Time-OrderedCameras] CNN 963 0 1.56 Asude [Aydin2023APerception] Hybrid 235 81 0.404 Asude [Aydin2023APerception] SNN 0.6 123 0.053 BDETT [Ding2022BiologicallyNetworks] SNN 0.7 115 0.046 Ours SNN 0.4 105 0.038